The Sorcerer

William Astle February 10, 2004

Notes

This work is copyright (c) 2004 by William Astle.

This work is licensed under the Creative Commons Attribution-NoDerivs-NonCommercial License. To view a copy of this license, visit http://creativecommons.org/licenses/by-nd-nc/1.0/ or send a letter to Creative Commons, 559 Nathan Abbott Way, Stanford, California 94305, USA.

I am also greatly interested in any feedback anyone may have. You can send such feedback by email to lost@l-w.net. You may also be able to find more works by me at my web site at http://lost.l-w.net/ so please stop by.

The Sorcerer

It was a bright summer day, the first of the season. Nor was enjoying the simple pleasure of sitting on his porch watching the forest and its denizens. As he watched the squirrels dashing here and there, he pondered his lot in life. He had had a longer life than anyone seemed entitled; indeed, he was, perhaps, the only one who remembered the lean times, let alone the collapse that precipitated them. Indeed, he even remembered the era of technology before the collapse. Fat lot of good all the technology did when the plague had come and wiped out almost all of humanity. Very few people knew how old Nor really was, and he liked it that way. It saved many awkward questions. Why was it that when someone found out you had lived for a century and a half they always felt they had to make some inane comment about what a gift it must be to see so much? Well, the older folks never did ask those sorts of questions; they simply looked at him with a knowing expression. They, after all, had sufficient life experience to realize what it might be like to live on. Still, he did have to admit to himself that his brand of continued life had a few things to recommend it. After all, he still looked and felt as though he were in his mid twenties rather than going on his second century. That fact alone probably saved him a lot of those awkward questions simply because nobody would have guessed his age from his appearance.

As he pondered the implications of very long life, he began to wonder, as he always did, why he had been gifted (or cursed depending his mood) with apparent immortality. No matter how much he pondered the question, he never did arrive at a satisfactory answer. This day, however, it occurred to him for the first time that he might not be the only one who was so gifted (or cursed). It followed that if he had managed to live for nearly two centuries, perhaps others had as well. Perhaps if he found others, there would be a common thread that would explain everything. But then he realized how unlikely it would be to find others like him; they could hardly be going around advertising their existence. If they had been, he would have no need to wonder if they existed. Indeed, they likely kept their existence secret for much the same reasons he did - undying people made other people nervous.

He supposed it was also possible that most of them had died already from one cause or another. For he had discovered long ago that even though he apparently did not age, nor did he fall ill, he still could be injured physically. While those injuries had had no lasting effects - no scars or permanent damage of any kind - the fact that he could be injured indicated that it must be possible to die. Though no expert on human anatomy, he was certain that he was as vulnerable as anyone to a sword through the heart or a bullet in the brain. However, he did seem to heal faster than most which would likely make some injuries that would kill others much less deadly. Indeed, it was likely, he mused, that secondary infections would be much less of a concern to him than they would for others; after all, whatever allowed his immune system to fight off any disease he had encountered to date would likely still work when he was injured.

The sun had climbed nearly to its zenith when his reverie was interrupted by a stranger walking up the path toward the cabin. This stranger was perhaps Nor’s height with long black hair. The stranger was also female unless his years of isolation had dimmed his eyesight and she had stunning blue eyes. As was typical of most folks during these times, she wore simple practical clothing with no attempt to emphasize her obvious attributes. She also had the air of having been travelling for a long time. As she approached his position on the porch, Nor found himself watching her intently. He had never had any visitors out this far, let alone a beautiful woman.

”It’s not polite to stare.” The statement shook Nor back to reality and he realized that she had covered the remaining ground to the porch.

”Sorry, I don’t get many visitors. I hardly remember how to behave among people,” Nor said, feeling somewhat silly for no reason he could put his finger on. ”What brings you out to the middle of nowhere?”

”I came to see for myself if the rumours were true.”

Nor was genuinely puzzled by this. ”Which rumours would those be?”

The visitor appeared somewhat amused by this. ”You mean to say that you have no idea what the folks in the village say about the strange hermit living deep in the forest?”

”Not the foggiest. I very seldom go anywhere near the village. Although I imagine they’ve come up with some rather inventive stories.” He allowed himself a very slight smile.

”No doubt you encouraged them somewhat in that endeavour.”

”Perhaps. Perhaps not. I expect you have a reason for coming all the way out here. I mean, other than the sharp wit.”

”I do. I wanted to meet the great Nor.”

Nor was unable to detect any trace of deceit but he was curious. ”And who is this Nor person?”

”You mean to say you are not him? All the evidence points here.” She seemed disappointed but Nor was not certain he bought it. Somehow he was sure that if she had come all this way to find him, her information must be very good.

”I believe that is what I implied. I am known as Hermit.” She raised her eyebrows at that. ”Not very creative, I know, but that is what I am known as.

What should I call you?”

”I have been called many things. Of late, the name that has been most predominant is Hunter.”

Nor and Hunter shared a brief chuckle of understanding. Then Nor said, ”Won’t you come in? You’ve clearly travelled a long way. Or perhaps you would prefer refreshments served on the porch?”

”The porch will be sufficient. The day is good and it would seem a shame to waste it.”

Nor disappeared into his cabin and reappeared moments later with two tall glasses of lemonade. Hunter seemed surprised by this but said nothing. The two of them sat on the porch and discussed various random subjects for the next several hours. Embedded in their conversation was a well fought match of verbal sparring during which neither party learned anything about the other that was not intended to be learned. By sunset, Nor’s last doubts about how Hunter had managed to find him were gone. She was as sharp as they came. But that still did not clarify why she had bothered to look for him anyway. History did record a man named Nor involved in the founding of Norstown shortly after the collapse yet that was long enough past that nobody would believe that that particular Nor could still be alive.

As the evening closed in, Nor kindled a fire in the pit in front of the cabin and the two of them sat watching the fire for a time. They talked of science and philosophy long into the night, their verbal sparring match long forgotten as pointless. Time seemed to stretch before them as the conversation droned on. The fire and the porch remained but everything else seemed to recede into unrealism. Nor realized that something significant was occurring but was not entirely certain what. What he did know was the Hunter had something to do with it. He wondered exactly how long she had been looking for him. It was now clear that she had intended to do whatever she was doing all along.

For lack of a better plan, Nor decided to observe as the situation developed. He watched as the unrealizing of the background continued. All sound beyond the sound of their own voices and the crackle of the fire had disappeared as had all other light sources. The air no longer seemed to move at all and the fire burned smokelessly without heat. Something was overwhelmingly familiar about this experience but he couldn’t quite bring that familiarity to the fore. As he glanced at his wrist chronometer, he observed that it, too, while still present had apparently ceased to function. It continued to display a time but that time was no longer changing. This only served to increase the feeling he should be remembering something. Realizing that he was not going to be able to force his memory to retrieve that information, he decided to focus on what was happening.

Now, he and Hunter were facing each other across the fire which was now burning in isolation. There was no ground, no sky, no feeling of heat or cold, all nature seemed to have departed. Realizing that his could not be the case, Nor concluded that they two had departed instead yet this did not explain what was happening. Still, the conversation continued between the two of them, as thought it had a life of its own. He was only dimly aware of it as he observed his current environment.

Suddenly, it hit him. He was not observing his surroundings with his eyes. Or his ears. Or touch. Or smell. As he realized this, the conversation became more immediate and he was aware of it as if it was the only thing happening. He realized that he must be exercising some sort of mental perception. He took a few minutes to become used to the dual sensations and was amazed how easy it was to differentiate them. Indeed, he began to realize that he had been exercising some level of this mental ability for a long time. Then he wondered that if he could divide his consciousness into two streams, could he add a third?

As he experimented with this mental ability, he discovered that it took no apparent effort to manage three streams of consciousness. Beyond that, it became difficult to keep things separate. Perhaps with practise he would be able to manage it. However, for the moment, three streams allowed him to continue the conversation with Hunter, ponder the situation, and observe his current situation. Even though he understood what he was doing with his mind, it still felt eerie.

He then turned his observer thread to analysing Hunter in the mind scape - for lack of a better term. He realized that there were slight differences in her appearance in the mind scape and on the porch. Her appearance was more idealized in the mind scape. He supposed that was probably the case with his appearance as well. As his eyes made contact with hers in the mind scape, however, he realized that he had made a potential mistake. He immediately felt a presence attempting to invade his mind. Without thinking, he raised a shield against the invasion. Strangely enough, the shield itself seemed to require no effort, as though it were a practised defence. He didn’t waste any time pondering that, however, as he diverted most of his mental resources from his pondering thread to the mind scape. As he did so, he remarked a slight indication of surprise in Hunter’s eyes, both in the mind scape and on the porch. Apparently he shouldn’t have been able to do that. He entertained a stray thought about how it was amazing what one could do when one didn’t know what one couldn’t do.

As he probed at the intrusion in the mind scape, he said on the porch, ”Out of curiosity, do you believe in magic?” and in the mind scape he said, ”Have you ever seen a wizard’s duel?”. Both of these statements he made simultaneously although he had no idea why he had made the latter one. After all, when could he have possibly witnessed a wizard’s duel. Even as Hunter strengthened her probe at his mind in the mind scape and attempted to answer his question on the porch, he suddenly realized why this experience had been familiar. With that figured out, he immediately directed all the mental resources he could conjure up toward the mind scape but left enough of a presence on the porch to handle whatever Hunter might do there. Then he went on the offensive.

Even as he strengthened his defences, he reached out with his mind toward Huntress. As he approached, he could see that she had split her consciousness into two threads, one for the porch and one for the mind scape. However, he detected just a bit of stress in her manner as he began to probe for weaknesses in her defences. Then he realized what she had said in response to his pair of questions. She had said, ”Rather obvious now, isn’t it?” on the porch and ”You have!?” in the mind scape. He wondered about that second answer but filed it for later. Now, he concentrated on battering her defences. In a short period of time, she began to sweat on the porch despite the chill night air and the fact that the fire had burned down (although the one in the mind scape was burning as brightly as ever). The strain began to show through in her voice as she attempted to continue the conversation. Then she gave up all pretense on the porch. Nor left a thin thread of his consciousness there in case anything happened. He suspected she had done the same and switched from inane chatter to a commentary on history.

Slowly he beat back at her defences. As he did so, her probes weakened and became less frequent. Very slowly, cracks (for lack of a better description) appeared in her defences. As he probed at those cracks, he felt her gathering strength. Immediately he slammed up a secondary defence, just as his instinctive shield shattered. Her probe slammed hard into his new shield and left him reeling mentally from the impact. However, the shield held and within a few seconds, both her attack and her defence evaporated. He had beaten her back. The last attack had left him slightly weakened but other than that, he felt fine.

Once he realized Hunter’s attack was thwarted, he shifted his attention primarily back to the porch. As he did so, he observed that the fire in the mind scape dissipated. Hunter’s presence shifted fully back to the porch in exhaustion and once he was certain of that, he let go his last thread in the mind scape.

”That was most impressive, Hermit. I have never encountered anyone with so much strength. Most, even with years of discipline, would not have defeated those attacks without showing signs of fatigue afterward. Yet you, without any training whatsoever, and hardly any understanding of what was going on, quite handily put me in my place. I thank you for doing nothing more than that.”

Nor accepted the compliment that was intended and said, ”You should sleep now. That fight took a lot out of you. If you’ll follow me, you can use my bed tonight. I have a lot of pondering to do and I doubt I’ll be needing it tonight.”

With Hunter comfortable in the cabin, Nor returned to the porch and switched on an electric light that he had managed to create during his tenure in the cabin. He wondered if he was being rash in trusting this Hunter after what she did. But then, he wondered if he would have believed her if she had just come out and told him what he had learned tonight. He might have intellectually, but he knew he would never have accepted it viscerally. It was clear that without that visceral understanding, self doubt would have crippled anything he tried to do. Perhaps by blind siding him with this she had done him a favour.

Then he remembered the feeling of familiarity he’d had during the experience. He thought about that for a moment. Perhaps this had not been his first experience with such things. Perhaps that had given him a leg up on Hunter. But that still didn’t explain the relative ease with which he had stopped her attacks, even that last desperate drive. Then he thought back to what he thought might have been his first experience with magic and realized what the similarity was.

He had been on vacation at the time, around the turn of the twenty first century, not long before the collapse. He had taken a trip to the Yukon to escape the rush and had been approaching the arctic circle when it happened. As he approached the final corner in the road before the parking lot for the marker, he remembered feeling a presence. At the time he’d been unable to pin it down but now that he thought about it, it felt like another mind. And that mind had been calling out for help, in desperation. Somehow he had responded. Even now he was not clear on what happened at that time, only that moments later, he remembered a strange experience as he was halfway around that corner. He remembered a strange temporal effect where he was still moving yet the various gauges and dials on his automobile’s dashboard and his chronometer had stopped moving. As had his car. Then he had seen a brilliant streak cross his field of view and shortly after everything had returned to normal leaving him feeling tired more than anything. He wondered if he had had anything to do with that streak across the sky.

With a start, he realized the sky was showing signs of dawn. He also realized that he was famished; he had had nothing to eat since the previous morning. He rose from his chair on the porch and began making preparations for breakfast. By the time the sun was fully in the sky, he had prepared a rather large breakfast (and an unhealthy one according to the experts but what did he care?). The aroma of the cooking had woken Hunter and the pair of them enjoyed the meal in silence. No words needed to be said as each understood the other perfectly at that moment.

Later that day, Hunter made preparations to leave. Just as she finished, she said, ”Welcome to the world of magic. Perhaps you’ll consider changing your name to ’Sorcerer’ now, rather than Hermit? Or perhaps you would consider coming with me?”

”While I appreciate what you did for me here, I have a need for solitude for the moment. Perhaps we will meet again at a later date,” Nor replied.

”Well, my mission was successful. I am satisfied that, whoever you truly are, you are not a threat to the continued existence of the human race. The very fact that you did nothing except win the fight, and you took no liberties with me in my weakened condition proves that. Had you killed me or otherwise rendered it impossible to return to my fellows, you would have been deemed hostile and hunted down. I am glad that will not be necessary, although I expect you would have been more than capable of defeating anyone they could have sent against you. You are possibly the strongest sorcerer in the history of this planet.” Hunter then shouldered her pack and set off back the way she came, exactly twenty four hours after she arrived.

Nor never expected to learn the name of the mysterious Hunter who taught him so much that day. Neither did he think Hunter would ever learn for sure that he truly was the ”great Nor” as she had put it. He was sure she suspected, but he was also absolutely certain she took that suspicion to her grave. He sincerely hoped the events of the previous twenty four hours had taught her as much as they had him. If so, she had the makings of a brilliant leader, and by all the powers, the Earth needed a few of them.

He switched off the light that he had forgotten was on and resumed his seat on the porch. As much as he tried to delude himself, however, life did not return to the old familiar pattern. Perhaps Hunter had done much more than she had set out to do after all. Perhaps she had saved him from himself. Only time would tell.